OpenStudy (nikeboy1996):

I will give a MEDAL! Do the data in the table represent a direct variation or an inverse variation? *data table posted after question

OpenStudy (carniel):

Sorry, we do not encourage user's to just do someone else's work. We only help other try to get the answer.

OpenStudy (nikeboy1996):

x 1 3 5 10 y 4 12 20 40

OpenStudy (nikeboy1996):

and Im not doing someone else's work, Im asking for help

OpenStudy (solidicecream):

Direct and inverse refer to how the two variables change compared to each other. Direct: When one increases, the other increases. When one decreases, the other decreases. Inverse (think opposite): When one increases the other decreases.

OpenStudy (nikeboy1996):

direct variation; y=4x <------ my answer direct variation; y = (1/4)x inverse variation; xy = 4 inverse variation; xy= 1/4

OpenStudy (nikeboy1996):

is my answer correct?

OpenStudy (solidicecream):

Hmmm Well I am not that good with math but I think you are right.

OpenStudy (solidicecream):

I mostly do biology sorry.

OpenStudy (carniel):

You said do the data in the table. Basically, telling us to do your work.

OpenStudy (nikeboy1996):

oops, I forgot to change the subject I was in, but thanks for helping me! :)

OpenStudy (carniel):

Also, your in the wrong section. Please go to Mathematics section to get help

OpenStudy (nikeboy1996):

ok, if you werent going to help, then y r u speaking?

OpenStudy (solidicecream):

Well no problem I am here to help!

OpenStudy (carniel):

I'm free to say what I want when ever and where ever. Also, if your doing something you aren't suppose to, I'm instructed to help.

OpenStudy (solidicecream):

He just wanted someone to make sure that his work was correct.

OpenStudy (carniel):

I was just telling him what he shouldn't be doing. No need to get mad over something.

OpenStudy (solidicecream):

I am not mad.... I just wanted to help that's why openstudy exist is to help people.

OpenStudy (carniel):

I know why it exists. I've been here for 2 years

OpenStudy (solidicecream):

I will not keep this argument going I am done here.

OpenStudy (nikeboy1996):

good for you! btw, Im a girl, my brother gave it to me bcuz he doesnt need it anymore bcuz he graduated high school, just so yk! Same thanks again@SolidIceCream ! :)

OpenStudy (carniel):

Just so I know? I'm graduating less than a month.